OpenStudy (anonymous):

write 4y-5x=3(4x-2y+1) in standard form

OpenStudy (anonymous):

multiply 3 4y-5x=12x-6y+3 add 5 x 4y = 17x - 6y + 3 subtract 4y 0 = 17x - 10y + 3

OpenStudy (anonymous):

thanks u deserve a medal!!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome